The president of Hill Enterprises, Terri Hill, projects the firms aggregate demand requirements over the next 8 months as follows: 1,400 1,700 1,800 1,900 2,300 2,200 1,900 1,900 May June January February March April August Her operations manager is considering a new plan, which begins in January with 200 units of inventory on hand. Stockout cost of lost sales is $125 per unit. Inventory holding cost is $20 per unit per month. Ignore any idle-time costs. The plan is called plan A.
Answer #1
Period Month Demand Production Ending(units) Stockout
0 Dec 1600 1600 200
1 Jan 1400 1600 400
2 Feb 1700 1400 100
3 Mar 1800 1700 0
4 Apr 1900 1800 100
5 May 2300 1900 300
6 June 2200 2300 100 0
7 July 1900 2200 300
8 Aug 1900 1900 300

Total stockout = 300 units

Total cost of stockout = 300 * 125 = $37,500

Inventory holding cost = (200 + 400 + 100 + 100 + 100 + 300 + 300 ) * 20 = $30,000

Period Month Demand Production Hire Fire
0 Dec 1600 1600 0 0
1 Jan 1400 1600 0 200
2 Feb 1700 1400 300 0
3 Mar 1800 1700 100 0
4 Apr 1900 1800 100 0
5 May 2300 1900 400 0
6 Jun 2200 2300 0 100
7 Jul 1900 2200 0 300
8 Aug 1900 1900 0 0

Overtime production cost =( 300 + 100 + 100 + 400 ) * 50 = $45000